my problem is
in any triangle the sum of the measures of the angles is 180 degrees. In triangle ABC, angle A is four times as large as angle B. Angle C measures 20 degrees less than angle B. Find the measure of each angle.
we are suposed to turn it into an equation find an angle then solve the problem from there. i cannot think of an equation.
i have tried:
180= 4X+X-20
and other equations but none work
thank you for helping me!

Answer by mananth(16946)

let angle B be x
Angle A = 4x
Angle C = x-20
..
x+4x+x-20 =180
6x-20=180
+20
6x=180+20
6x=200
x= 33.33 Angle B
..
Angle A = 4x = 133.32
Angle C =x-20= 13.33
CHECK
33.33 + 133.32 + 13.33